# Discrete Distribution Networks

### v1

@@ Abstract
We introduce a novel generative model, the Discrete Distribution Networks (DDN), that approximates data distribution using hierarchical discrete distributions. We posit that since the features within a network inherently contain distributional information, liberating the network from a single output to concurrently generate multiple samples proves to be highly effective. Therefore, DDN fits the target distribution, including continuous ones, by generating multiple discrete sample points. To capture finer details of the target data, DDN selects the output that is closest to the Ground Truth (GT) from the coarse results generated in the first layer. This selected output is then fed back into the network as a condition for the second layer, thereby generating new outputs more similar to the GT. As the number of DDN layers increases, the representational space of the outputs expands exponentially, and the generated samples become increasingly similar to the GT. This hierarchical output pattern of discrete distributions endows DDN with two intriguing properties: highly compressed representation and more general zero-shot conditional generation. We demonstrate the efficacy of DDN and these intriguing properties through experiments on CIFAR-10 and FFHQ.
1 Introduction
@@
Figure 2: DDN enables zero-shot conditional generation. DDN also supports conditions in non-pixel domains, such as text-to-image with CLIP [22]. Images enclosed in yellow borders serve as the ground truth. The abbreviations in the table header correspond to their respective tasks as follows: ‘SR’ stands for Super-Resolution, with the following digit indicating the resolution of the condition. ‘ST’ denotes Style Transfer, which computes Perceptual Losses with the condition according to [11].
@@
With the advent of ChatGPT [3] and DDPM [9], deep generative models have become increasingly popular and significant in everyday life. However, modeling the complex and diverse high-dimensional data distributions is challenging. Previous methods [21, 2, 6, 28, 16, 14, 27, 7] have each demonstrated their unique strengths and characteristics in modeling these distributions. Most methods, apart from autoregressive models [29], have introduced complex concepts or operations [1]. To address this, we propose Discrete Distribution Networks as a simpler, more direct form of generative modeling.
@@
Most image generation models applied in daily life are conditional generative models. They generate corresponding images based on content provided by users, such as images to be edited , reference images [25], text descriptions, hand-drawn editing strokes [33], sketches, and so on. Current mainstream generative models [34, 24, 23] require training separate models and parameters for each condition. Each model has a fixed format for conditions and lacks an easy way to adjust the influence of each condition, limiting users’ creative possibilities. DDN support a wide range of zero-shot conditional generation, bringing more possibilities for user creativity. Fig. 2 demonstrates the capabilities of DDN in zero-shot conditional generation.
@@
The core concept of Discrete Distribution Networks (DDN) is to approximate the distribution of training data using a multitude of discrete sample points. The secret to generating diverse samples lies in the network’s ability to concurrently generate multiple samples ($K$). This is perceived as the network outputting a discrete distribution – the origin of the name Discrete Distribution Networks. All generated samples serve as the sample space for this discrete distribution. Typically, each sample in this discrete distribution has an equal probability mass of $1/K$. Our goal is to make this discrete distribution as close as possible to the target dataset.
@@
To accurately fit the target distribution of large datasets, a substantial representational space is required. In the most extreme scenario, this space must be larger than the number of training data samples. However, current neural networks lack the feasibility to generate such a vast number of samples simultaneously. Therefore, we adopt a strategy from autoregressive models [31] and partition this large space into a hierarchical conditional probability model. Each layer of this model needs only a small number of outputs. We then select one of these outputs as the output for that layer and use it as conditional input to the next layer. As a result, the output of the next layer will be more closely related to the selected conditional sample. If the number of layers is $L$ and the number of outputs per layer is $K$, then the output space of the network is $K^{L}$. Due to its exponential nature, this output space will be much larger than the number of samples in the dataset. Fig. 1 shows how DDN generates images.
We posit that the contributions of this paper are as follows:
  • We introduce a novel generative model, termed Discrete Distribution Networks (DDN), which exhibit a more straightforward and streamlined principle and form.
  • For training the DDN, we propose the “Split-and-Prune” optimization algorithm, and a range of practical techniques.
  • We conduct preliminary experiments and analysis on the DDN, showcasing its intriguing properties and capabilities, such as zero-shot conditional generation and highly compact representations.
@@ 2 Related Work
Figure 3: Schematic of Discrete Distribution Networks (DDN). (a) The data flow during the training phase of DDN is shown at the top. As the network depth increases, the generated images become increasingly similar to the training images. Within each Discrete Distribution Layer (DDL), $K$ samples are generated, and the one closest to the training sample is selected as the generated image for loss computation. These $K$ output nodes are optimized using Adam with the Split-and-Prune method. The right two figures demonstrate the two model paradigms supported by DDN. (b) Single Shot Generator Paradigm: Each neural network layer and DDL has independent weights. (c) Recurrence Iteration Paradigm: All neural network layers and DDLs share weights. For inference, replacing the Guided Sampler in the DDL with a random choice enables the generation of new images.
Deep Generative Model. Generative Adversarial Networks (GANs) [21, 2] and Variational Autoencoders (VAEs) [16, 32] are two early successful generative models. GANs reduce the divergence between the generated sample distribution and the target distribution through a game between the generator and discriminator. However, regular GANs cannot map samples back to the latent space, thus they cannot reconstruct samples. VAEs encode data into a simple distribution’s latent space through an Encoder, and the Decoder is trained to reconstruct the original data from this simple distribution’s latent space. Notably, VQ-VAE [32], an important improvement of VAEs, achieves a more compact discrete representation by replacing the VAE’s latent space with discrete codebooks. However, this discrete representation still has a 2D structure, leading to a significant amount of redundant information. Moreover, VQ-VAE relies on a prior network to perform generative tasks. Autoregressive models [29] , with their simple principles and methods, model the target distribution by decomposing the target data into conditional probability distributions of each component. They can also compute the exact likelihood of target samples. However, the efficiency of these models is reduced when dealing with image data, which is not suitable to be decomposed into a sequence of components. Normalizing flow [14] is another class of generative models that can compute the likelihood. They use invertible networks to construct a mapping from samples to a noise space, and during the generation stage, they map back from noise to samples. Energy Based Models (EBM) [6, 28] with their high-quality and rich generative results, have led to the rise of diffusion models. However, their multi-step iterative generation process requires substantial computational resources.
Conditional image generation. Early conditional image generation methods were primarily dominated by GAN-based methods [10, 36, 19]. However, these methods require a considerable amount of paired condition-image data and necessitate training different models for distinct condition tasks. Recent works have demonstrated the excellent zero-shot conditional generation capacity of diffusion models [19, 18]. One approach is to add a moderate amount of noise to the image-based condition, ensuring that the condition’s structure is not completely destroyed by the noise, and making it quite similar to an image containing noise at a certain step in the diffusion training process. The denoising process of diffusion then yields realistic images. However, these diffusion-based zero-shot conditional generation methods cannot handle non-image-based conditions, such as generating samples of a specific category under the guidance of a classifier.
@@ 3 Discrete Distribution Networks
Figure 4: Illustration of the Split-and-Prune operation. For example in (a), the light blue bell-shaped curve represents a one-dimensional target distribution. The 5 “” under the x-axis are the initial values from a uniform distribution of 5 output nodes, which divide the entire space into 5 parts using midpoints between adjacent nodes as boundaries (i.e., vertical gray dashed lines). Each part corresponds to the range represented by this output node on the continuous space $x$. Below each node are three values: $P$ stands for the relative frequency of the ground truth falling within this node’s range during training; $Q$ refers to the probability mass of this sample (node) in the discrete distribution output by the model during the generation phase, which is generally equal for each sample, i.e., $1/K$. The bottom-most value denotes the difference between $P$ and $Q$. Colorful horizontal line segments represent the average probability density of $P$, $Q$ within corresponding intervals. In (b), the Split operation selects the node with the highest $P$ (circled in red). In (c), the Prune operation selects the node with the smallest $P$ (circled in red). In (d), through the combined effects of loss and Split-and-Prune operations, the distribution of output nodes moves towards final optimization. From the observed results, the KL divergence ($KL(P||Q)$) consistently decreases as the operation progresses, and the yellow line increasingly approximates the light blue target distribution.
3.1 Network Architecture
Fig. 2(a) illustrates the overall structure of the DDN, comprised of Neural Network Blocks and Discrete Distribution Layers (DDL). Each DDL contains $K$ output nodes, each of which is a set of 1x1 convolutions responsible for transforming the feature into the corresponding output image. The parameters of these 1x1 convolutions are optimized by Adam with Split-and-Prune. The $K$ images generated by the $K$ output nodes are inputted into the Guided Sampler. The Guided Sampler selects the image with the smallest L2 distance to the training image, which serves as the output of the current layer and is used to calculate the L2 loss with the training image. Simultaneously, the selected image is concatenated back into the feature, acting as the condition for the next block. The index (depicted in green as K-1 in Fig. 2(a)) of the selected image represents the latent value of the training sample at this layer. Through the guidance of the Guided Sampler layer by layer, the image generated by the network progressively becomes more similar to the training sample until the final layer produces an approximation of the training sample. We average the L2 loss across each layer as the final loss of the entire network.
When the network performs the generation task, replacing the Guided Sampler with a random choice enables image generation. Given the exponential representational space of $K^{L}$ sample space and the limited number of samples in the training set, the probability of sampling an image with the same latent space as those in the training set is also exponentially low. For image reconstruction tasks, the process is almost identical to the training process, only substituting the training image with the target image to be reconstructed and omitting the L2 Loss part from the training process. The Final Generated Image in Fig. 2(a) represents the final reconstruction result. The indices of the selected samples along the way form the latent of the target image. Therefore, the DDN latent is a sequence of integers with length $L$, which we regard as the hierarchical discrete representation of the target sample. The latent space exhibits a tree structure with $L$ layers and $K$ degrees per node, where each leaf node represents a sample space, and its latent representation indicates the indices of all nodes along the path to this leaf node, as shown in Fig. 0(b). In the latent sequence, values placed earlier correspond to higher-level nodes in the tree, controlling the low-frequency information of the output sample, while later values tend to affect high-frequency information.
Figure 5: Toy examples for two-dimensional data generation. The numerical values at the bottom of each figure represent the Kullback-Leibler (KL) divergence. Due to phenomena such as “dead nodes” and “density shift”, the application of Gradient Descent alone fails to properly fit the Ground Truth (GT) density. However, by employing the Split-and-Prune strategy, the KL divergence is reduced to even lower than that of the Real Samples. More instructions for this experiment are in the Appendix.
3.2 Optimiztion with Split-and-Prune
@@
Algorithm 1 Split-and-Prune for Training
1:Output nodes number $K$, model $f$, non-output parameters $\boldsymbol{\theta}$, target distribution $q(\mathbf{x})$
2:Initialize output node parameters $\mathbf{\boldsymbol{\psi}}(k)$ for $k\in\{1,\dotsc,K\}$ with random values
@@
[…]
9:    Gradient descent $\nabla_{\boldsymbol{\theta},\boldsymbol{\psi}(k^{*})}\left\|f(\boldsymbol{\theta},\boldsymbol{\psi}(k^{*}))-\mathbf{x}\right\|^{2}$
@@
15:         $\mathbf{\boldsymbol{\psi}}(k_{new}):=\operatorname{clone}(\mathbf{\boldsymbol{\psi}}(k_{max}))$
@@
21:         $\operatorname{remove}\;\mathbf{\boldsymbol{\psi}}(k_{min})$ and $\mathbf{c}(k_{min})$

### v2

@@ Abstract
We introduce a novel generative model, the Discrete Distribution Networks (DDN), that approximates data distribution using hierarchical discrete distributions. We posit that since the features within a network inherently capture distributional information, enabling the network to generate multiple samples simultaneously, rather than a single output, may offer an effective way to represent distributions. Therefore, DDN fits the target distribution, including continuous ones, by generating multiple discrete sample points. To capture finer details of the target data, DDN selects the output that is closest to the Ground Truth (GT) from the coarse results generated in the first layer. This selected output is then fed back into the network as a condition for the second layer, thereby generating new outputs more similar to the GT. As the number of DDN layers increases, the representational space of the outputs expands exponentially, and the generated samples become increasingly similar to the GT. This hierarchical output pattern of discrete distributions endows DDN with unique property: more general zero-shot conditional generation. We demonstrate the efficacy of DDN and its intriguing properties through experiments on CIFAR-10 and FFHQ. The code is available at https://discrete-distribution-networks.github.io/
@@ 1 Introduction
Figure 2: DDN enables more general zero-shot conditional generation. DDN supports zero-shot conditional generation across non-pixel domains, and notably, without relying on gradient, such as text-to-image generation using a black-box CLIP model Radford et al. (2021). Images enclosed in yellow borders serve as the ground truth. The abbreviations in the table header correspond to their respective tasks as follows: ‘SR’ stands for Super-Resolution, with the following digit indicating the resolution of the condition. ‘ST’ denotes Style Transfer, which computes Perceptual Losses with the condition according to Johnson et al. (2016).
@@
With the advent of ChatGPT Brown et al. (2020) and DDPM Ho et al. (2020), deep generative models have become increasingly popular and significant in everyday life. However, modeling the complex and diverse high-dimensional data distributions is challenging. Previous methods Kingma & Welling (2014); Radford et al. (2016); Kingma & Dhariwal (2018); Goyal et al. (2021); Song et al. (2021); Shocher et al. (2024); Graves et al. (2023) have each demonstrated their unique strengths and characteristics in modeling these distributions. In this work, we propose a novel approach to model the target distribution, where the core idea is to generate multiple samples simultaneously, allowing the network to directly output an approximate discrete distribution. Hence, we name our method Discrete Distribution Networks (DDN). DDN embraces a core concept as simple as autoregressive models, offering another straightforward and effective form for generative models.
@@
Most generative models applied in real-world scenarios are conditional generative models. Taking image generation as an example, these models generate corresponding images based on content provided by users, such as images to be edited, reference images Saharia et al. (2021), text descriptions, hand-drawn editing strokes Voynov et al. (2022), sketches, and so on. Current mainstream generative models Rombach et al. (2021); Ramesh et al. (2022); Zhang et al. (2023) typically require training separate models and parameters for each condition. These models are restricted to fixed condition formats and lack the flexibility to adjust the influence of each condition dynamically, thereby limiting users’ creative freedom.
Recent works have attempted to address this issue through zero-shot conditional generation (ZSCG). However, these methods either only support conditions in the same pixel domain as the training data Wang et al. (2022); Lugmayr et al. (2022); Meng et al. (2021); Nair et al. (2023) or depend on discriminative models to supply gradients during generation Yu et al. (2023). In contrast, DDN supports a wide range of ZSCG tasks, encompassing both pixel-domain and non-pixel-domain conditions, as shown in fig. 2. To the best of our knowledge, DDN is the first generative model capable of performing zero-shot conditional generation in non-pixel domains without relying on gradient information. This implies that DDN can achieve ZSCG solely based on black-box discriminative models.
@@
The core concept of Discrete Distribution Networks (DDN) is to approximate the distribution of training data using a multitude of discrete sample points. The secret to generating diverse samples lies in the network’s ability to concurrently generate multiple samples ($K$). This is perceived as the network outputting a discrete distribution. All generated samples serve as the sample space for this discrete distribution. Typically, each sample in this discrete distribution has an equal probability mass of $1/K$. Our goal is to make this discrete distribution as close as possible to the target dataset.
@@
To accurately fit the target distribution of large datasets, a substantial representational space is required. In the most extreme scenario, this space must be larger than the number of training data samples. However, current neural networks lack the feasibility to generate such a vast number of samples simultaneously. Therefore, we adopt a strategy from autoregressive models Van Den Oord et al. (2016) and partition this large space into a hierarchical conditional probability model. Each layer of this model needs only a small number of outputs. We then select one of these outputs as the output for that layer and use it as conditional input to the next layer. As a result, the output of the next layer will be more closely related to the selected conditional sample. If the number of layers is $L$ and the number of outputs per layer is $K$, then the output space of the network is $K^{L}$. Due to its exponential nature, this output space will be much larger than the number of samples in the dataset. fig. 1 shows how DDN generates images.
We posit that the contributions of this paper are as follows: (1) We introduce a novel generative model, termed Discrete Distribution Networks (DDN), which exhibit a more straightforward and streamlined principle and form. (2) For training the DDN, we propose the “Split-and-Prune” optimization algorithm, and a range of practical techniques. (3) We conduct preliminary experiments and analysis on the DDN, showcasing its intriguing properties and capabilities, such as zero-shot conditional generation and highly compact representations.
@@ 2 Related Work
Deep Generative Model. Generative Adversarial Networks (GANs) Radford et al. (2016); Brock et al. (2019) and Variational Autoencoders (VAEs) Kingma & Welling (2014); van den Oord et al. (2018) are two early successful generative models. GANs reduce the divergence between the generated sample distribution and the target distribution through a game between the generator and discriminator. However, regular GANs cannot map samples back to the latent space, thus they cannot reconstruct samples. VAEs encode data into a simple distribution’s latent space through an Encoder, and the Decoder is trained to reconstruct the original data from this simple distribution’s latent space. Notably, VQ-VAE van den Oord et al. (2018), an important improvement of VAEs, achieves a more compact discrete representation by replacing the VAE’s latent space with discrete codebooks. However, this discrete representation still has a 2D structure, leading to a significant amount of redundant information. Moreover, VQ-VAE relies on a prior network to perform generative tasks. Autoregressive models van den Oord et al. (2016) , with their simple principles and methods, model the target distribution by decomposing the target data into conditional probability distributions of each component. They can also compute the exact likelihood of target samples. However, the efficiency of these models is reduced when dealing with image data, which is not suitable to be decomposed into a sequence of components. Normalizing flow Kingma & Dhariwal (2018) is another class of generative models that can compute the likelihood. They use invertible networks to construct a mapping from samples to a noise space, and during the generation stage, they map back from noise to samples. Energy Based Models (EBM) Goyal et al. (2021); Song et al. (2021) with their high-quality and rich generative results, have led to the rise of diffusion models. However, their multi-step iterative generation process requires substantial computational resources. The recent Idempotent Generative Network Shocher et al. (2024) introduces a novel approach by training a neural network to be idempotent, mapping any input to the target distribution effectively.
@@
Figure 3: Schematic of Discrete Distribution Networks (DDN). (a) The data flow during the training phase of DDN is shown at the top. As the network depth increases, the generated images become increasingly similar to the training images. Within each Discrete Distribution Layer (DDL), $K$ samples are generated, and the one closest to the training sample is selected as the generated image for loss computation. These $K$ output nodes are optimized using Adam with the Split-and-Prune method. The right two figures shown the two model paradigms supported by DDN. (b) Single Shot Generator Paradigm: Each neural network layer and DDL has independent weights. (c) Recurrence Iteration Paradigm: All neural network layers and DDLs share weights. For inference, replacing the Guided Sampler in the DDL with a random choice enables the generation of new images.
@@ 3 Discrete Distribution Networks
Network architecture. Figure 3(a) illustrates the overall structure of the DDN, comprised of Neural Network Blocks and Discrete Distribution Layers (DDL). Each DDL contains $K$ output nodes, each of which is a set of 1x1 convolutions responsible for transforming the feature into the corresponding output image. The parameters of these 1x1 convolutions are optimized by Adam with Split-and-Prune. The $K$ images generated by the $K$ output nodes are inputted into the Guided Sampler. The Guided Sampler selects the image with the smallest L2 distance to the training image, which serves as the output of the current layer and is used to calculate the L2 loss with the training image. Simultaneously, the selected image is concatenated back into the feature, acting as the condition for the next block. The index (depicted in green as “K-1” in fig. 3(a)) of the selected image represents the latent value of the training sample at this layer. Through the guidance of the Guided Sampler layer by layer, the image generated by the network progressively becomes more similar to the training sample until the final layer produces an approximation of the training sample.
Objective function. The DDN model consists of $L$ layers of Discrete Distribution Layers (DDL). For a given layer $l$, denoted as $f_{l}$, the input is the selected sample from the previous layer, $\mathbf{x}^{*}_{l-1}$. The layer generates $K$ new samples, $f_{l}(\mathbf{x}^{*}_{l-1})$, from which we select the sample $\mathbf{x}^{*}_{l}$ that is closest to the current training sample $\mathbf{x}$, along with its corresponding index $k_{l}^{*}$. The loss $J_{l}$ for this layer is then computed only on the selected sample $\mathbf{x}^{*}_{l}$.
$k_{l}^{*}=\underset{k\in\{1,\dots,K\}}{\operatorname{argmin}}\;\left\|f_{l}(\mathbf{x}^{*}_{l-1})[k]-\mathbf{x}\right\|^{2}$ (1)
$\mathbf{x}^{*}_{l}=f_{l}(\mathbf{x}^{*}_{l-1})[k_{l}^{*}]\quad;\quad J_{l}=\left\|\mathbf{x}^{*}_{l}-\mathbf{x}\right\|^{2}$ (2)
Here, $\mathbf{x}^{*}_{0}=\mathbf{0}$ represents the initial input to the first layer. For simplicity, we omit the details of input/output feature, neural network blocks and transformation operations in the equations.
By recursively unfolding the above equations, we can derive the latent variable $\mathbf{k}^{*}_{1:L}$ and the global objective function $J$.
$\mathbf{k}^{*}_{1:L}=\left[k_{1}^{*},k_{2}^{*},\dots,k_{L}^{*}\right]=\left[\underset{k\in\{1,\dots,K\}}{\operatorname{argmin}}\;\left\|\mathcal{F}([\mathbf{k}^{*}_{1:l-1},k])-\mathbf{x}\right\|^{2}\right]_{l=1}^{L}$ (3)
$J=\frac{1}{L}\sum_{l=1}^{L}\left\|\mathcal{F}(\mathbf{k}^{*}_{1:l})-\mathbf{x}\right\|^{2}$ (4)
Here, $\mathcal{F}$ represents the composite function formed from $f_{l}$, defined as: $\mathcal{F}(\mathbf{k}_{1:l})=f_{l}(f_{l-1}(\dots f_{1}(\mathbf{x}_{0})[k_{1}]\dots)[k_{l-1}])[k_{l}]$. Finally, we average the L2 loss across all layers to obtain the final loss for the entire network.
How to generation. When the network performs the generation task, replacing the Guided Sampler with a random choice enables image generation. Given the exponential representational space of $K^{L}$ sample space and the limited number of samples in the training set, the probability of sampling an image with the same latent space as those in the training set is also exponentially low. For image reconstruction tasks, the process is almost identical to the training process, only substituting the training image with the target image to be reconstructed and omitting the L2 Loss part from the training process. The Final Generated Image in fig. 3(a) represents the final reconstruction result. The indices of the selected samples along the way form the latent representation of the target image $\mathbf{k}^{*}_{1:L}$, same as eq. 3. Therefore, the latent $\mathbf{k}^{*}_{1:L}$ is a sequence of integers with length $L$, which we regard as the hierarchical discrete representation of the target sample. The latent space exhibits a tree structure with $L$ layers and $K$ degrees per node, where each leaf node represents a sample space, and its latent representation indicates the indices of all nodes along the path to this leaf node, as shown in fig. 1(b). In the latent sequence, values placed earlier correspond to higher-level nodes in the tree, controlling the low-frequency information of the output sample, while later values tend to affect high-frequency information.
@@
Figure 4: Illustration of the principle behind the Split-and-Prune operation. For example in (a), the light blue bell-shaped curve represents a one-dimensional target distribution. The 5 “” under the x-axis are the initial values from a uniform distribution of 5 output nodes, which divide the entire space into 5 parts using midpoints between adjacent nodes as boundaries (i.e., vertical gray dashed lines). Each part corresponds to the range represented by this output node on the continuous space $x$. Below each node are three values: $P$ stands for the relative frequency of the ground truth falling within this node’s range during training; $Q$ refers to the probability mass of this sample (node) in the discrete distribution output by the model during the generation phase, which is generally equal for each sample, i.e., $1/K$. The bottom-most value denotes the difference between $P$ and $Q$. Colorful horizontal line segments represent the average probability density of $P$, $Q$ within corresponding intervals. In (b), the Split operation selects the node with the highest $P$ (circled in red). In (c), the Prune operation selects the node with the smallest $P$ (circled in red). In (d), through the combined effects of loss and Split-and-Prune operations, the distribution of output nodes moves towards final optimization. From the observed results, the KL divergence ($KL(P||Q)$) consistently decreases as the operation progresses, and the yellow line increasingly approximates the light blue target distribution.
3.1 Optimization with Split-and-Prune
@@
Figure 5: Toy examples for two-dimensional data generation. The numerical values at the bottom of each figure represent the Kullback-Leibler (KL) divergence. Due to phenomena such as “dead nodes” and “density shift”, the application of Gradient Descent alone fails to properly fit the Ground Truth (GT) density. However, by employing the Split-and-Prune strategy, the KL divergence is reduced to even lower than that of the Real Samples. More details for this experiment are in the Appendix.
Algorithm 1 Split-and-Prune of one layer
1:Output nodes number $K$, model $f$, non-output parameters $\bm{\theta}$, target distribution $q(\mathbf{x})$
2:Initialize output node parameters $\mathbf{\bm{\psi}}(k)$ for $k\in\{1,\dotsc,K\}$ with random values
@@
8:     Choose $k^{*}=\underset{k\in\bm{\psi}}{\operatorname{argmin}}\;\left\|f(\bm{\theta},\bm{\psi}(k))-\mathbf{x}\right\|^{2}$
9:     Gradient descent $\nabla_{\bm{\theta},\bm{\psi}(k^{*})}\left\|f(\bm{\theta},\bm{\psi}(k^{*}))-\mathbf{x}\right\|^{2}$
@@
15:         $\mathbf{\bm{\psi}}(k_{new}):=\operatorname{clone}(\mathbf{\bm{\psi}}(k_{max}))$
@@
21:         $\operatorname{remove}\;\mathbf{\bm{\psi}}(k_{min})$ and $\mathbf{c}(k_{min})$